Fisk corporation is trying to improve its inventory control system and has installed an online computer at its retail stores. Fisk anticipates sales of 75,000 units per year, an ordering cost of $8 per order, and carrying costs of $1.20 per until.
a) what is the economic ordering quantity?
b)how many orders will be placed during the year?
c) what will the average inventory be?
d) what is the total cost of ordering and carrying inventory?
Answer #1

a. economic ordering quantity = square root of (2*demand for the year*cost to place a single order/cost to hold one unit inventory for a year)

= square root of (2*75,000*8)/1.20 = 1000 units

b. How many orders = 75,000/1,000 = 75 orders

c. average inventory = eoq/2 = 1000/2 = 500 units

d. ordering cost = 75*8 = $600
carrying cost = 500*1.20 = $600
total $1200